# PO Box 129188, Saadiyat Island, Abu Dhabi, United Arab Emirates

### v1

@@ 1 Introduction
In general, there are huge uncertainties on the history of the Universe before the Big Bang nucleosynthesis epoch Allahverdi:2020bys ; Batell:2024dsi . In particular, the dynamics of cosmic reheating, that is, on how the energy stored in the inflaton is transferred to the SM thermal bath, is largely unknown. During the reheating phase, the Hubble expansion of the Universe was dominated by the inflaton. Its energy density is generally considered to evolve like non-relativistic matter or radiation, which are associated with an inflaton oscillating at the bottom of a quadratic or quartic potential, respectively. However, it can also scale faster than radiation, as in the case of kination Spokoiny:1993kt ; Ferreira:1997hj , or even more faster, as in the context of ekpyrotic Khoury:2001wf ; Khoury:2003rt or cyclic models Gasperini:2002bn ; Erickson:2003zm ; Barrow:2010rx ; Ijjas:2019pyf , with further discussion in Ref. Scherrer:2022nnz . Moreover, the inflaton could decay or annihilate into various types of SM particles, leading to different dependencies of the temperature of the SM bath on the cosmic scale factor. In particular, during reheating, the SM temperature could increase, decrease, or even stay constant Giudice:2000ex ; Co:2020xaf .
In this work, we demonstrate that the phenomenology of the DM UV freeze-in production strongly depends on the dynamics of the cosmic reheating era. Using a general parametrization for the Hubble expansion rate and SM temperature, we carefully study the DM production during reheating. This allows us not only to recover previous results reported in the literature, but also to systematically explore large unexplored regions of the parameter space of reheating. In particular, we study the DM UV freeze-in production from decays of the inflaton, and annihilations parametrized by general temperature dependence. Notably, the most general parameterization for inflaton-decay scenarios has not been addressed in prior studies. As motivated examples, we specifically studied the cases of gravitational DM production from the scattering of SM particles or inflatons Ema:2015dka ; Garny:2015sjg ; Tang:2016vch ; Ema:2016hlw ; Tang:2017hvq ; Garny:2017kha ; Bernal:2018qlk ; Ema:2018ucl ; Almeida:2018oid ; Ema:2019yrd ; Chianese:2020yjo ; Bernal:2020bfj ; Ahmed:2020fhc ; Bernal:2020qyu ; Chianese:2020khl ; Kolb:2020fwh ; Bernal:2020ili ; Redi:2020ffc ; Ling:2021zlj ; Mambrini:2021zpp ; Bernal:2021kaj ; Barman:2021ugy ; Barman:2021qds ; Haque:2021mab ; Clery:2021bwz ; Barman:2022tzk ; Clery:2022wib ; Garcia:2022vwm ; Kaneta:2022gug ; Basso:2022tpd ; Barman:2022qgt ; Haque:2023yra ; Kolb:2023dzp ; Garcia:2023awt ; Garcia:2023qab ; Kaneta:2023uwi ; Barman:2023opy ; Garcia:2023obw ; Kolb:2023ydq ; Racco:2024aac ; Dorsch:2024nan ; Choi:2024bdn ; Verner:2024agh ; Jenks:2024fiu ; Lee:2024wes ; Henrich:2024rux ; Xu:2024cey .
@@ 2 Parametrizing Cosmic Reheating
Figure 1: Summary of the different reheating scenarios. The black dot corresponds to the standard case where the inflaton scales as non-relativistic matter and decays into SM particles with a constant decay width, while the black crosses correspond to the alternative scenarios described in the text. The red area in the upper left corner does not give rise to viable reheating. The central blue dashed line corresponds to $k_{c}=8$; above, over and below it, the gravitational DM production from SM scatterings has an power-law, logarithmic or order-one enhancement during reheating, respectively. For reference, the lines $k_{c}=6$ and $k_{c}=12$ are also shown.
@@ 3 UV Dark Matter Production
where $\Lambda$ corresponds to an effective energy scale of new physics, and $k>5$ being a constant. Concrete examples of this rate will be presented in Section 4.
@@ 3.1 After Reheating
Finally, we emphasize that, in the present scenario, possible thermalization and number-changing processes within the dark sector are ignored. They can have a strong impact, in particular by enhancing the DM relic abundance by several orders of magnitude Bernal:2020gzm . For the sake of simplicity, we assume here that there are no sizable self-interactions within the dark sector.
@@ 3.2.1 Through Scatterings
Several comments are in order: $i)$ This result reproduces the behavior found in Ref. Garcia:2017tuj for a massive inflaton decaying into SM particles with a constant decay width (that is, $\omega=0$ and $\alpha=3/8$), and in Ref. Bernal:2019mhf , for general equation-of-state parameters $\omega$ and $\alpha=\frac{3}{8}(1+\omega)$. $ii)$ The overall prefactor of Eq. (25) has the same scaling shown in Eq. (21), however, new factors depending on $k$ appear. $iii)$ For $\alpha\,k<\alpha\,k_{c}$ an order-one boost occurs. $iv)$ In the case where $k=k_{c}$ the enhancement is logarithmic and larger than one, interdependently of the sign of $\alpha$. $v)$ A potentially large power-law boost can take place only if $\alpha\,k>\alpha\,k_{c}$. It occurs only for $\alpha>0$, since $T_{I}>T_{\text{rh}}$ is required. $vi)$ We note that Eq. (25) behaves well in the limit $\alpha=0$.
@@ 4.2 Inflaton scatterings
[…]

### v2

@@ 1 Introduction
In general, there are huge uncertainties on the history of the Universe before the Big Bang nucleosynthesis epoch Allahverdi:2020bys ; Batell:2024dsi . In particular, the dynamics of cosmic reheating, that is, on how the energy stored in the inflaton is transferred to the SM thermal bath, is largely unknown Hasegawa:2019jsa . During the reheating phase, the Hubble expansion of the Universe was dominated by the inflaton. Its energy density is generally considered to evolve like non-relativistic matter or radiation, which are associated with an inflaton oscillating at the bottom of a quadratic or quartic potential, respectively. However, it can also scale faster than radiation, as in the case of kination Spokoiny:1993kt ; Ferreira:1997hj , or even more faster, as in the context of ekpyrotic Khoury:2001wf ; Khoury:2003rt or cyclic models Gasperini:2002bn ; Erickson:2003zm ; Barrow:2010rx ; Ijjas:2019pyf , with further discussion in Ref. Scherrer:2022nnz . Moreover, the inflaton could decay or annihilate into various types of SM particles, leading to different dependencies of the temperature of the SM bath on the cosmic scale factor. In particular, during reheating, the SM temperature could increase, decrease, or even stay constant Giudice:2000ex ; Co:2020xaf .
In this work, we demonstrate that the phenomenology of the DM UV freeze-in production strongly depends on the dynamics of the cosmic reheating era. Using a general parametrization for the Hubble expansion rate and SM temperature, we carefully study the DM production during reheating. This allows us not only to recover previous results reported in the literature, but also to systematically explore large unexplored regions of the parameter space of reheating. In particular, we study the DM UV freeze-in production from decays of the inflaton, and annihilations parametrized by general temperature dependence. Notably, the most general parameterization for inflaton-decay scenarios has not been addressed in prior studies. As motivated examples, we specifically studied the cases of gravitational DM production from the scattering of SM particles or inflatons Ema:2015dka ; Garny:2015sjg ; Tang:2016vch ; Ema:2016hlw ; Tang:2017hvq ; Garny:2017kha ; Bernal:2018qlk ; Ema:2018ucl ; Almeida:2018oid ; Ema:2019yrd ; Chianese:2020yjo ; Bernal:2020bfj ; Ahmed:2020fhc ; Bernal:2020qyu ; Chianese:2020khl ; Kolb:2020fwh ; Bernal:2020ili ; Redi:2020ffc ; Ling:2021zlj ; Mambrini:2021zpp ; Bernal:2021kaj ; Barman:2021ugy ; Barman:2021qds ; Haque:2021mab ; Clery:2021bwz ; Barman:2022tzk ; Clery:2022wib ; Garcia:2022vwm ; Kaneta:2022gug ; Basso:2022tpd ; Barman:2022qgt ; Haque:2023yra ; Kolb:2023dzp ; Garcia:2023awt ; Garcia:2023qab ; Kaneta:2023uwi ; Barman:2023opy ; Garcia:2023obw ; Kolb:2023ydq ; Racco:2024aac ; Dorsch:2024nan ; Choi:2024bdn ; Verner:2024agh ; Jenks:2024fiu ; Lee:2024wes ; Henrich:2024rux ; Xu:2024cey ; Maleknejad:2024hoz .
@@ 2 Parametrizing Cosmic Reheating
Figure 1: Summary of the different reheating scenarios. The black dot corresponds to the standard case where the inflaton scales as non-relativistic matter and decays into SM particles with a constant decay width, while the black crosses correspond to the alternative scenarios described in the text. The red area in the upper left corner does not give rise to viable reheating. The central blue dashed line corresponds to $k_{c}\equiv\frac{3}{2}\frac{\omega+3}{\alpha}=8$; above, over and below it, the gravitational DM production from SM scatterings has an power-law, logarithmic or order-one enhancement during reheating, respectively. For reference, the lines $k_{c}=6$ and $k_{c}=12$ are also shown.
@@ 3 UV Dark Matter Production
where $k>5$ being a constant and $\Lambda$ corresponds to an effective energy scale of new physics. For this approach to be valid, $\Lambda$ must be much higher than the temperatures reached by the SM thermal bath and the DM mass. Concrete examples of the rate $\gamma_{a}$ will be presented in Section 4.
@@ 3.1 After Reheating
Finally, we emphasize that, in the present scenario, possible thermalization and number-changing processes within the dark sector are ignored. They can have a strong impact, in particular by enhancing or even reducing the DM relic abundance by several orders of magnitude Bernal:2020gzm ; Bhattiprolu:2023akk ; Bhattiprolu:2024dmh . For the sake of simplicity, we assume here that there are no sizable self-interactions within the dark sector.
@@
Before closing this section, we emphasize that the previous discussion only applies to the production of DM particles lighter than the reheating temperature $m<T_{\text{rh}}$. Heavier DM particles with mass $m>T_{\text{rh}}$ could also be produced after reheating, but their production rate density in Eq. (12) must be multiplied by a Boltzmann suppression factor $\exp[-m/T]$, reflecting the fact that only the tail of the SM thermal distribution contributes to the production of DM; see, e.g. Refs. Cosme:2023xpa ; Koivunen:2024vhr ; Arcadi:2024wwg ; Arcadi:2024obp . This approach corresponds to an instantaneous reheating. In the next section, instead, we study the genesis of DM during a non-vanishing reheating phase.
@@ 3.2.1 Through Scatterings
Several comments are in order: $i)$ This result reproduces the behavior found in Ref. Garcia:2017tuj for a massive inflaton decaying into SM particles with a constant decay width (that is, $\omega=0$ and $\alpha=3/8$), and in Ref. Bernal:2019mhf , for general equation-of-state parameters $\omega$ and $\alpha=\frac{3}{8}(1+\omega)$. $ii)$ The overall prefactor of Eq. (25) has the same scaling shown in Eq. (21), however, new factors depending on $k$ appear. $iii)$ For $\alpha\,k<\alpha\,k_{c}$ an order one boost could occur, as the temperature dependence is identical. However, the ratio of DM production during and after reheating $\frac{g_{\star s}(T_{0})}{g_{\star s}(T_{\text{rh}})}\frac{k-5}{(k_{c}-k)\,\alpha}$ is typically smaller than one, making DM production during reheating subdominant. $iv)$ In the case where $k=k_{c}$ the enhancement is logarithmic and larger than one, interdependently of the sign of $\alpha$. $v)$ A potentially large power-law boost can take place only if $\alpha\,k>\alpha\,k_{c}$. It occurs only for $\alpha>0$, since $T_{I}>T_{\text{rh}}$ is required. $vi)$ We note that Eq. (25) behaves well in the limit $\alpha=0$.
@@ 4.2 Inflaton scatterings
$\Lambda=T_{\text{rh}}\left(\frac{64\pi}{9}\,\frac{{m_{\phi}^{\text{rh}}}^{2}\,T_{\text{rh}}^{4}}{m^{2}\,H_{\text{rh}}^{4}}\right)^{\frac{\alpha}{2(3-2\alpha)}},$ (35)
and $k=6/\alpha$. As in this case where $k$ is no longer constant and depends on $\alpha$, the equivalent of the critical exponent $k_{c}$ becomes a critical equation-of-state parameter $\omega_{c}=1$. Even if the general expressions in Section 3.2.1 could be easily translated to this case, for the sake of clarity, we prefer to explicitly report how Eq. (25) reflects to this case:
$Y_{\text{rh}}\simeq\frac{1}{128}\,\sqrt{\frac{g_{\star}}{10}}\,\frac{g_{\star}}{g_{\star s}}\frac{T_{\text{rh}}^{3}\,m^{2}}{M_{P}^{3}\,{m_{\phi}^{\text{rh}}}^{2}}\times\begin{dcases}\frac{1}{1-\omega}\left(\frac{T_{I}}{T_{\text{rh}}}\right)^{\frac{3}{2}\frac{1-\omega}{\alpha}}&\text{ for }\omega<1\,,\\ \frac{3}{2\alpha}\,\ln\frac{T_{I}}{T_{\text{rh}}}&\text{ for }\omega=1\,,\\ \frac{1}{\omega-1}&\text{ for }\omega>1\,.\end{dcases}$ (36)
The left panel of Fig. 6 shows the parameter space required to fit the entire DM abundance through gravitational annihilation of inflatons for $m=10^{12}$ GeV, ${m_{\phi}^{\text{rh}}}=10^{13}$ GeV, and $\omega=0$, taking $|\alpha|=3/4$ (blue), $|\alpha|=9/16$ (green), $|\alpha|=3/8$ (black) and $\alpha=0$ (black dot). It only takes into account the contribution during reheating, where inflatons dominate the energy density of the Universe. The positive and negative values for $\alpha$ are depicted with solid and dashed lines, respectively. The red dotted line corresponds to $T_{I}=T_{\text{rh}}$. The yellow areas are excluded because the temperature is higher than the inflaton mass, while the red areas correspond to the CMB constraint on the inflationary scale for the three values of $\alpha$. This choice for the equation-of-state parameter is described by the first case of Eq. (36) (e.g. $\omega<1$) and therefore a strong dependence on the ratio $T_{I}/T_{\text{rh}}$ is expected, as can be observed in the figure. The different lines converge when $T_{I}=T_{\text{rh}}$, corresponding to an instantaneous reheating era, and to the case $\alpha=0$ where the temperature is constant during reheating. For $\alpha>0$, very large values of $T_{I}$ are excluded due to the constraint on the inflationary scale. However, for $\alpha<0$, $T_{I}<T_{\text{rh}}$ and much smaller values of $T_{I}$ can give rise to the abundance of DM observed. Interestingly, and in contrast to the previous case of scattering of SM particles, here both positive and negative values for $\alpha$ give rise to a power-law enhancement.
Figure 6: Parameter space required to fit the entire DM abundance through gravitational annihilation of inflatons for $m=10^{12}$ GeV, ${m_{\phi}^{\text{rh}}}=10^{13}$ GeV, $\omega=0$ (left) or $\omega=1/3$ (right), and different values of $\alpha$, for fermionic DM. The red dotted lines correspond to $T_{I}=T_{\text{rh}}$. The yellow areas are excluded because the temperature is higher than the inflaton mass. The red areas correspond to the the combination of the constraints from CMB and the inflaton mass, for the three values of $\alpha$.
The right panel of Fig. 6 compares to the left panel, but for $\omega=1/3$, taking $|\alpha|=1$ (blue), $|\alpha|=5/8$ (green) and $|\alpha|=1/4$ (black). For this case, the parameter space that reproduces the observed DM abundance is generally similar; however, the constraint on the inflaton mass becomes more involved, since $m_{\phi}$ is no longer constant. The red bands correspond to the combination of the constraints from the CMB and the inflaton mass. For $T_{I}>T_{\text{rh}}$ a stronger upper bound appears, coming from Eq. (17), while for $T_{I}<T_{\text{rh}}$ a lower bound from Eq. (18) exists for the three values of $\alpha$.
5 Conclusions
In this work, we have studied the phenomenology of dark matter (DM) production in the early Universe through the ultraviolet (UV) freeze-in (FIMP) paradigm. In this scenario, the bulk of the DM is generated at the highest temperatures reached by the standard model (SM) thermal plasma, typically during the cosmic reheating era. Taking into account that the dynamics of reheating is largely unknown, we used a general parametrization for the Hubble expansion rate and SM temperature as a function of the cosmic scale factor, to bracket the uncertainty of the reheating epoch. This approach allows to recover standard reheating scenarios such as the cases where $i)$ the inflaton oscillates at the bottom of a monomial potential while decaying or annihilating into SM states, $ii)$ the inflaton decays with a general field-dependent decay width, $iii)$ the SM temperature during reheating is constant, $iv)$ or even scenarios where the energy density inflaton of the inflaton gets diluted faster than radiation so that its decay is not required. However, we emphasize that this approach also allows the exploration of a wide variety of alternative reheating histories, as long as the Hubble expansion and the SM temperature behave as simple power laws.
Within this approach to reheating, we analytically computed the production of DM in the early Universe from decays and annihilations, using a general temperature dependence for the interaction rates. Depending on the details of the reheating era and the DM production rate, the DM production could peak at the beginning or the end of reheating, and the final DM relic abundance could have a logarithmic or a power-law dependence on the ratio of temperatures at the beginning and the end of reheating. We explored the parameter space required to match the observed DM abundance, bracketing systematic uncertainties related to the reheating dynamics. Previous results reported in the literature were recovered, while viable alternative reheating evolutions were also investigated. This generic setup was used in the context of two particularly well-motivated examples: The gravitational DM production from the scattering of SM particles and scatterings of inflatons.
Before closing, we note that the present scenario is challenging to test experimentally for several reasons. $i)$ Due of its very suppressed interactions with the visible sector, the UV freeze-in paradigm evades detection with standard techniques. $ii)$ The observables sensitive to the cosmic reheating period are scarce. However, novel detectors such as the proposed Windchime experiment Windchime:2022whs are sensitive to heavy DM, even if it only interacts gravitationally with the SM. Additionally, gravitational waves provide a promising method for probing the early Universe and, in particular, the reheating era Ghiglieri:2015nfa ; Ghiglieri:2020mhm ; Ringwald:2020ist ; Muia:2023wru ; Villa:2024jkw ; Frey:2024jqy ; Villa:2024jbf ; Bernal:2024jim ; Xu:2024cey . As we look forward, we are excited to see progress in our theoretical grasp of cosmic reheating, which can enhance our comprehension of the early Universe and offer fresh perspectives on the nature of DM.
Acknowledgements.
The authors thank Sasha Pukhov for helpful conversations and Yong Xu for comments on the draft. NB received funding from the Grant PID2023-151418NB-I00 funded by MCIU/AEI/10.13039/501100011033/ FEDER, UE.